OpenStudy (anonymous):

The sum of a geometric series is -2044. If the first term is -4 and the common ratio is 2, what is the final term in the sequence?

OpenStudy (anonymous):

Sn = a[r^n-1]/r-1

OpenStudy (anonymous):

use this and find n

OpenStudy (anonymous):

Wait @Yahoo!

OpenStudy (anonymous):

then an = a*r^(n-1)

OpenStudy (anonymous):

Why @waterineyes any mistake

OpenStudy (anonymous):

Sorry my mistake... Ha ha ha..

OpenStudy (anonymous):

\[-2044 = -4[2^n -1]/(2-1)\]

OpenStudy (anonymous):

511 = 2^n - 1 510 = 2^n

OpenStudy (anonymous):

find n??

OpenStudy (rsadhvika):

\(-2044 = -4. \frac{2^{n}-1}{2-1}\) \(-2044 = -4. (2^{n}-1)\) \(511 = 2^{n}-1\) \(512 = 2^{n}\) \(2^9 = 2^{n}\) 9=n

OpenStudy (anonymous):

512

OpenStudy (anonymous):

sorry 512 = 2^n

OpenStudy (anonymous):

now n = 9

OpenStudy (anonymous):

subs in an = a * r^(n-1)

OpenStudy (anonymous):

\[\large 2^n = 512 \implies 2^n = 2^9 \implies \color{green}{n = 9}\]

OpenStudy (anonymous):

an = -4 * 2^8

OpenStudy (anonymous):

you can replace n here by 9 @Yahoo!

OpenStudy (anonymous):

256 * -4 = -1024

OpenStudy (anonymous):

lol thanks!

OpenStudy (anonymous):

Welcome